# A Comprehensive Survey on Retrieval Methods in Recommender Systems

### v1

@@ 1. INTRODUCTION
Figure 2. Taxonomy for retrieval methods in recommender systems.
@@ 1.1. A Taxonomy and Major Content
We review numerous works published in major conferences and journals and summarize the efforts made to achieve precise and efficient personalized retrieval into three main areas, as is illustrated in Figure 2:
@@
The remaining of this article is structured as follows: Section 2 introduces the background of retrieval in recommender systems, including problem formulation, basic retrieval strategies and input data for retrieval. Sections 3 and  4 detail the efforts that focus on similarity learning using shallow and deep structures, respectively. Section 5 discusses improvements in indexing mechanisms for efficient retrieval. Section 6 covers optimization methods of retrieval, including key aspects of model learning, such as widely-used loss functions and different negative sampling strategies. Section 7 summarizes commonly used datasets and conducts detailed experiments on three public datasets to assess the effectiveness of various methods. Section 8 presents a specific case study of current industrial practices in the retrieval stage at a specific company. We discuss open challenges and suggest promising future research directions in Section 9, and conclude the survey in Section 10.
@@ 2.2.2. Personalized Retrieval
Personalized retrieval recommends items that align with user interests. While popular items may draw broad attention, they do not always match individual user preferences. Hence, personalized retrieval aims to cater to users’ dynamic information needs and maintain their stickiness. As is shown in Figure  3 (b), U2I, U2I2I (often shortened as I2I for simplicity), and U2U2I are three popular personalized retrieval strategies, where ‘U’ stands for the user, ‘I’ stands for the item, and ‘2’ is short for ‘to’ that denotes the process of seeking similarity or connection. For instance, U2I directly correlates the target user with items he/she might like; U2I2I identifies items similar to those the target user has previously interacted with; and U2U2I finds users with similar tastes to suggest items for the target user. Each path introduced above offers a unique strategy to discover and recommend items that user might be interested in.
@@ 2.3.2. Rich Side Information
In addition to the user-item rating matrix, three types of side information play crucial roles in enhancing retrieval methods in recommender systems (Shi et al., 2014), as illustrated in Figure  3 (d). One vital source is the user profile, which might include demographics information, such as a user’s gender, age, and hobbies. Item attributes also provide rich information, detailing properties of the items like their category or content. Contextual information provides insights into the circumstances under which user interactions with items occur (Adomavicius and Tuzhilin, 2010). Commonly included in this category are timestamps, which track when interactions like ratings or purchases take place (Xiong et al., 2010; Koren, 2009). Additionally, other relevant details like a user’s location during an app download (Böhmer et al., 2011) or their hunger status when rating a food menu (Ono et al., 2009) can also enrich the understanding of user-item interactions, further tailoring the retrieval process to specific user conditions or preferences.
@@ 2.4. Multi-Channel Retrieval
The choice of retrieval channel is highly dependent on the specific applications. For example, the ‘trending list’ and ‘interest tags’ are essential for timely platforms like news streaming, while ‘friend favorites’ is more important for social media platforms. The number of items to be retrieved (i.e., top-$K$) can vary for different channels, and would be typically optimized through offline evaluation and online A/B testing.
@@ 4.1. Two-Tower Methods
The two-tower methods represent a state-of-the-art approach in the retrieval stage, offering a balance between prediction accuracy and inference efficiency (Li et al., 2022). In these methods, the data used for training consists of instances denoted as $(x,y)$, where $y\in\{0,1\}$ indicates user feedback (Richardson et al., 2007). A label of 1 signifies positive feedback, and 0 denotes negative. The feature $x$ can be divided into user features and item features. In the two-tower architecture, these user and item features are processed independently in parallel sub-towers, namely, the user tower and the item tower. Taking the user tower as an example, user features are first passed through an embedding layer, where user feature embeddings are obtained via an embedding look-up operation (Zhang et al., 2016). These embeddings are then processed through multiple fully connected layers to refine the representation. The final output of the two-tower model is determined by the inner product of the user and item representations. This architecture decouples the user and item modeling, which is highly efficient and significantly reduces inference latency and computational resources. Classic implementations of two-tower methods in industry include Microsoft’s DSSM (Huang et al., 2013), Google’s YouTubeDNN (Covington et al., 2016), and Airbnb’s personalized embeddings (Grbovic and Cheng, 2018), leading to significant performance improvements. The two-tower methods mentioned above, while efficient, often face limitations in adequately modeling user-item interactions, which can impede model performance. To overcome such challenges, advanced techniques have been developed to enhance the depth of interaction between the user and item side, which we will further detail in future directions in Section 9.
@@ 4.2. Autoencoder Based Methods
Autoencoder based methods, particularly Variational Autoencoders (VAEs) (Kingma and Welling, 2013), are increasingly utilized in the retrieval stage of recommender systems to address challenges like data sparsity and the complexity of user-item interactions. These methods excel in generating and representing complex data structures. VAEs operate by assuming that observable data points are generated from latent variables following some probability distribution, such as standard Gaussian distribution. They consist of two main components: an encoder and a decoder. The encoder, or inference model, approximates the true posterior distribution of the latent variables given the input data, effectively capturing the underlying data structure. The decoder, on the other hand, reconstructs the input data from the latent representation, aiming to match the original input as closely as possible.
@@ 4.3. Graph Embedding Based Methods
Graph embedding based methods have been pivotal in enhancing the retrieval process in recommender systems. These methods extend the principles of word embeddings, such as Word2Vec (Mikolov et al., 2013), to the domain of recommendations. Item2Vec (Barkan and Koenigstein, 2016) adapts Word2Vec to items, treating the user behavior sequences of items (like products or songs) as sentences to uncover similarities based on co-occurrence. This enables better recommendations by capturing contextual similarities among items. For example, Airbnb successfully applied Item2Vec (Grbovic and Cheng, 2018) to personalize user experiences in online accommodations. DeepWalk (Perozzi et al., 2014) uses random walks on graphs to generate node sequences, learning representations that reflect the social structure of interactions. Node2Vec (Grover and Leskovec, 2016) enhances this approach by balancing breadth-first and depth-first sampling, capturing diverse connectivity patterns in the graph. Further refining this, EGES (Wang et al., 2018a) integrates side information (like user demographics or item attributes) into the embeddings, enriching the model’s contextual understanding. LINE (Tang et al., 2015) preserves both direct connections and neighbor similarities, ensuring embeddings reflect both explicit and implicit relationships. SDNE (Wang et al., 2016) uses autoencoders to learn node representations, adeptly capturing nonlinear relationships within complex, hierarchical structures. All these models underscore the importance of graph embeddings in understanding and leveraging complex user-item interaction patterns.
@@ 4.4. Graph Neural Network Based Methods
Recent advancements have incorporated Graph Neural Network (GNN) techniques into the retrieval stage of recommender systems (Berg et al., 2017; He et al., 2020; Sun et al., 2019; Liu et al., 2020b), known for their efficacy in handling complex relational data. To understand the distinction between graph embedding based methods and GNN based methods, it is essential to first explain their foundational differences. Graph embedding based methods typically involve deriving item sequences from the graph and then generating embeddings from these sequences. These methods focus on capturing the structural properties and proximities within the graph to represent items in a continuous vector space. In contrast, GNN based methods utilize graph convolution operations, which involve iterative message passing and aggregation across the graph’s nodes and edges. GNNs excel in capturing intricate user-item interactions, optimizing the retrieval process by exploiting high-order relationships. Figure 4 illustrates the pipeline for applying GNNs in retrieval.
@@
GNN based retrieval methods show remarkable results in understanding and leveraging complex user-item interaction patterns. However, enhancing their efficiency for large-scale retrieval applications remains a critical research problem.
@@ 4.5. MLP Based Collaborative Filtering Methods
Unlike conventional collaborative filtering techniques that often rely on linear models and explicit factorization methods, neural collaborative filtering (NCF) (He et al., 2017) and its variants leverage the expressive power of multi-layer perceptron (MLP) to capture complex user-item interactions through non-linear optimization.  Bai et al. (2017) enhance NCF by incorporating local neighborhood information using convolutional filters, improving interaction representation between users and items.  Liu et al. (2017) propose PhD, a hybrid recommendation model combining user-item interactions with side information using probabilistic matrix factorization.  Cao et al. (2018) develop a group recommendation method integrating an attention mechanism with MLP within the NCF framework.  Chen et al. (2018b) utilize heterogeneous information networks (HIN) to represent diverse data types, capturing complex semantics through meta paths.  Fu et al. (2018) investigate both local and global co-occurrences in user-user and item-item relations to capture contextual information more comprehensively.  Zhou et al. (2019) enrich the model’s understanding by incorporating textual representations from items, enhancing both explicit and implicit feedback.  Chen et al. (2019a) combine deep feature learning and deep interaction modeling using two MLP networks with a rating matrix.  Otunba et al. (2019) propose an ensemble framework integrating generalized matrix factorization (GMF) and MLP networks, leveraging the strengths of deep and shallow architectures.  Xue et al. (2019) focus on capturing high-order item relations in item based CF models using an NCF based approach.  Yi et al. (2019a) develop deep matrix factorization (DMF), integrating diverse side information from users and items to enhance the retrieval process.  Chen et al. (2020) propose efficient whole-data learning strategies to avoid common sampling techniques in CF retrieval methods. Additionally, closely related methods like CML (Hsieh et al., 2017) is based on metric learning, and NBPO (Yu and Qin, 2020) employs noisy-label robust learning techniques for retrieval.

### v2

@@ Abstract.
Recommender System, Retrieval Stage, Candidate Generation
@@ 1.1. A Taxonomy and Major Content
Figure 2. Distributions of the reviewed papers of retrieval methods in recommender systems over years (a) and over venues (b).
We collect over 100 research papers on retrieval methods in recommender systems. In conducting our literature search, we utilize DBLP and Google Scholar as the primary search engines, using keywords such as candidate matching, candidate retrieval, large-scale recommendation, and collaborative filtering in the titles and abstracts to filter relevant papers from top-tier conferences and journals in data mining, machine learning, recommender systems, and artificial intelligence. These include, but are not limited to, KDD, ICDM, WWW, SIGIR, ICML, NIPS, AAAI, WSDM, CIKM, RecSys, TKDE, and TOIS, covering the period from 2001 to 2024, as shown in Figure 2. We traverse the citation graph of the identified papers and incorporate pertinent studies. In addition to reviewing published papers, we also screen preprints on arXiv, identifying those with novel and intriguing ideas to provide a more comprehensive perspective. Based on this corpus, we summarize the efforts made to achieve precise and efficient personalized retrieval into three main areas, as illustrated in Figure 3:
Figure 3. Taxonomy for retrieval methods in recommender systems.
@@
The remaining of this article is structured as follows: Section 2 introduces the background of retrieval in recommender systems, including problem formulation, basic retrieval strategies and input data for retrieval. Sections 3 and  4 detail the efforts that focus on similarity learning using shallow and deep structures, respectively. Section 5 discusses improvements in indexing mechanisms for efficient retrieval. Section 6 covers optimization methods of retrieval, including key aspects of model learning, such as widely-used loss functions and different negative sampling strategies. Section 7 presents commonly used evaluation metrics for both offline and online settings. Section 8 summarizes commonly used datasets and conducts detailed experiments on three public datasets to assess the effectiveness of various methods. Section 9 presents a specific case study of current industrial practices in the retrieval stage at a specific company. We discuss open challenges and suggest promising future research directions in Section 10, and conclude the survey in Section 11.
@@ 2.2.2. Personalized Retrieval
Personalized retrieval recommends items that align with user interests. While popular items may draw broad attention, they do not always match individual user preferences. Hence, personalized retrieval aims to cater to users’ dynamic information needs and maintain their stickiness. As is shown in Figure  4 (b), U2I, U2I2I (often shortened as I2I for simplicity), and U2U2I are three popular personalized retrieval strategies, where ‘U’ stands for the user, ‘I’ stands for the item, and ‘2’ is short for ‘to’ that denotes the process of seeking similarity or connection. For instance, U2I directly correlates the target user with items he/she might like; U2I2I identifies items similar to those the target user has previously interacted with; and U2U2I finds users with similar tastes to suggest items for the target user. Each path introduced above offers a unique strategy to discover and recommend items that user might be interested in.
@@ 2.3.2. Rich Side Information
In addition to the user-item rating matrix, three types of side information play crucial roles in enhancing retrieval methods in recommender systems (Shi et al., 2014), as illustrated in Figure  4 (d). One vital source is the user profile, which might include demographics information, such as a user’s gender, age, and hobbies. Item attributes also provide rich information, detailing properties of the items like their category or content. Contextual information provides insights into the circumstances under which user interactions with items occur (Adomavicius and Tuzhilin, 2010). Commonly included in this category are timestamps, which track when interactions like ratings or purchases take place (Xiong et al., 2010; Koren, 2009). Additionally, other relevant details like a user’s location during an app download (Böhmer et al., 2011) or their hunger status when rating a food menu (Ono et al., 2009) can also enrich the understanding of user-item interactions, further tailoring the retrieval process to specific user conditions or preferences.
@@
Empirical evidence indicates that incorporating rich side information yields substantial improvements over using only the user–item rating matrix. For instance, Tso-Sutter et al. (2008) report that fusing tag information with ratings increases Recall@10 from 0.32 to 0.34 compared to pure user and item based collaborative filtering on the Last.fm (son, 2024) dataset. SocialMF (Jamali and Ester, 2010) incorporates social trust propagation into matrix factorization, reducing RMSE from 1.175 to 1.075 on the Epinions dataset (epi, 2003) and from 0.878 to 0.821 on Flixster (fli, 2009) relative to basic matrix factorization. With respect to contextual data, Koren (2009) demonstrate that adding temporal dynamics reduces RMSE from 0.913 to 0.897 on the Netflix movie-rating dataset (Bennett and Lanning, 2007) compared to a static SVD++ baseline. Details on datasets appear in Section 8, and metric definitions appear in Section 7.
@@ 2.4. Multi-Channel Retrieval
The choice of retrieval channel is highly dependent on the specific applications. For example, the ‘trending list’ and ‘interest tags’ are essential for timely platforms like news streaming, while ‘friend favorites’ is more important for social media platforms. The number of items to be retrieved (i.e., top-$K$) can vary for different channels, and would be typically optimized through offline evaluation and online A/B testing (Huang et al., 2025).
2.5. Comparison with Multi-Stage Retrieval in Ad Hoc Search
Multi-stage cascaded pipelines are also well established in ad hoc search (Clarke et al., 2016; Chen et al., 2017). Clarke et al. (2016) introduce a framework that quantifies efficiency–effectiveness trade-offs in multi-stage query systems without relying on relevance judgments. Chen et al. (2017) propose a cost-aware cascade ranking model that jointly optimizes per-stage cutoffs to balance latency against retrieval quality in a static document index.
Although both ad hoc search and recommender system pipelines employ the same core principle, i.e., progressively narrowing a large corpus through increasingly expensive stages, they differ along several key dimensions:
  • Personalization and dynamics. Ad hoc search typically serves identical results to all users over a fixed index, whereas recommender systems have to tailor candidate sets to individual user profiles and adapt continuously as new items and interactions arrive.
  • Data freshness and indexing. Ad hoc indices are updated at regular intervals, making feature extraction relatively stable; recommender pipelines often require real-time or near-real-time indexing to incorporate the latest user behavior and item metadata.
  • Feature representations. Early ad hoc stages use term-based statistics (e.g., BM25 (Robertson et al., 1995)), while the retrieval stage in recommendation leverages learned embeddings that combine user history, content attributes, and contextual signals.
@@ 4.1. Two-Tower Methods
The two-tower methods represent a state-of-the-art approach in the retrieval stage, offering a balance between prediction accuracy and inference efficiency (Li et al., 2022). In this architecture, each user $u$ and item $i$ is independently mapped to a representation via two parallel sub-networks: $\phi_{u}(u;\theta_{u})$ and $\phi_{i}(i;\theta_{i})$, where $\phi_{u}$ and $\phi_{i}$ denote the user tower and item tower, respectively, and $\theta_{u},\theta_{i}$ are their parameters. Taking the user tower as an example, user features are first passed through an embedding layer, where user feature embeddings are obtained via an embedding look-up operation (Zhang et al., 2016). These embeddings are then processed through multiple fully connected layers to refine the representation. The final output of the two-tower model is determined by the inner product of the user and item representations.
(8) $s(u,i)=\left\langle\phi_{u}(u;\theta_{u}),\phi_{i}(i;\theta_{i})\right\rangle.$
This architecture decouples the user and item modeling, which is highly efficient and significantly reduces inference latency and computational resources. Classic implementations of two-tower methods in industry include Microsoft’s DSSM (Huang et al., 2013), Google’s YouTubeDNN (Covington et al., 2016), and Airbnb’s personalized embeddings (Grbovic and Cheng, 2018), leading to significant performance improvements. The two-tower methods mentioned above, while efficient, often face limitations in adequately modeling user-item interactions, which can impede model performance. To overcome such challenges, advanced techniques have been developed to enhance the depth of interaction between the user and item side, which we will further detail in future directions in Section 10.
@@ 4.2. Autoencoder Based Methods
Autoencoder based methods, particularly Variational Autoencoders (VAEs) (Kingma and Welling, 2013), are increasingly utilized in the retrieval stage of recommender systems to address challenges like data sparsity and the complexity of user-item interactions. These methods excel in generating and representing complex data structures. VAEs operate by assuming that each observed interaction vector $\mathbf{r}_{u}\in\{0,1\}^{N}$ is generated from a latent variable $\mathbf{z}$ following some probability distribution, such as standard Gaussian distribution. They consist of two main components: an encoder and a decoder. The encoder network $q_{\phi}(\mathbf{z}\mid\mathbf{r}_{u})$ approximates the true posterior distribution of the latent variables given the input data, effectively capturing the underlying data structure:
(9) $\bm{\mu}_{u}=\mu(\mathbf{r}_{u}),\quad\log\bm{\sigma}_{u}^{2}=\log\sigma^{2}(\mathbf{r}_{u}),$
(10) $\mathbf{z}_{u}=\bm{\mu}_{u}+\bm{\sigma}_{u}\odot\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$
The decoder then reconstructs $\mathbf{r}_{u}$ from the latent representation, as shown in  equation 11, where $\sigma(\cdot)$ denotes the element-wise sigmoid function. It aims to match the original input as closely as possible.
(11) $\hat{\mathbf{r}}_{u}=\sigma\bigl(\mathrm{Decoder}(\mathbf{z}_{u})\bigr),$
@@ 4.3. Graph Embedding Based Methods
Graph embedding based methods have been pivotal in enhancing the retrieval process in recommender systems. These methods extend the principles of word embeddings, such as Word2Vec (Mikolov et al., 2013), to user–item interaction graphs. These methods learn a mapping that assigns each node $v$ in the graph $G=(\mathcal{V},\mathcal{E})$ to a vector $\Phi(v)$. They optimize a neighborhood‐preserving objective of the form:
(12) $\min_{\Phi}\sum_{v\in\mathcal{V}}\sum_{c\in\mathcal{N}(v)}-\log\Pr\bigl(c\,|\,\Phi(v)\bigr),\quad\Pr\bigl(c|\Phi(v)\bigr)=\frac{\exp\bigl(\Phi(c)^{\top}\Phi(v)\bigr)}{\sum_{u\in\mathcal{V}}\exp\bigl(\Phi(u)^{\top}\Phi(v)\bigr)},$
where $\mathcal{N}(v)$ denotes a chosen context set for node $v$ (e.g., adjacent nodes, nodes co‐occurring in random walks, or nodes sharing side‐information). Minimizing this loss ensures that frequently co‐occurring or closely related nodes receive similar embeddings, while unrelated nodes are pushed apart. Item2Vec (Barkan and Koenigstein, 2016) is a direct adaptation of Word2Vec to items, treating the user behavior sequences of items (like products or songs) as sentences to uncover similarities based on co-occurrence. This enables better recommendations by capturing contextual similarities among items. For example, Airbnb successfully applied Item2Vec (Grbovic and Cheng, 2018) to personalize user experiences in online accommodations. DeepWalk (Perozzi et al., 2014) uses random walks on graphs to generate node sequences, learning representations that reflect the social structure of interactions. Node2Vec (Grover and Leskovec, 2016) enhances this approach by balancing breadth-first and depth-first sampling, capturing diverse connectivity patterns in the graph. Further refining this, EGES (Wang et al., 2018a) integrates side information (like user demographics or item attributes) into the embeddings, enriching the model’s contextual understanding. LINE (Tang et al., 2015) preserves both direct connections and neighbor similarities, ensuring embeddings reflect both explicit and implicit relationships. SDNE (Wang et al., 2016) uses autoencoders to learn node representations, adeptly capturing nonlinear relationships within complex, hierarchical structures. All these models underscore the importance of graph embeddings in understanding and leveraging complex user-item interaction patterns.
@@ 4.4. Graph Neural Network Based Methods
Recent advancements have incorporated Graph Neural Network (GNN) techniques into the retrieval stage of recommender systems (Berg et al., 2017; He et al., 2020; Sun et al., 2019; Liu et al., 2020b), known for their efficacy in handling complex relational data. To understand the distinction between graph embedding based methods and GNN based methods, it is essential to first explain their foundational differences. Graph embedding based methods typically involve deriving item sequences from the graph and then generating embeddings from these sequences. These methods focus on capturing the structural properties and proximities within the graph to represent items in a continuous vector space. In contrast, GNN based methods utilize graph convolution operations, which involve iterative message passing and aggregation across the graph’s nodes and edges. In a typical message-passing paradigm, for each node $v$ at layer $k$, one computes
(13) $m_{v}^{(k)}$ $=\mathrm{AGG}\bigl(\{\,h_{u}^{(k-1)}:u\in\mathcal{N}(v)\}\bigr),$
$h_{v}^{(k)}$ $=\mathrm{UPDATE}\bigl(h_{v}^{(k-1)},\,m_{v}^{(k)}\bigr),$
where $\mathcal{N}(v)$ is the set of neighbors of $v$, $h_{v}^{(k)}$ is the representation of node $v$ at layer $k$, $\mathrm{AGG}(\cdot)$ defines a pooling or aggregation function over neighbors (e.g., mean, sum, or attention), and $\mathrm{UPDATE}(\cdot)$ is a function—often a linear transformation followed by a nonlinearity—that integrates the previous representation with the aggregated message. GNNs excel at capturing intricate user-item interactions, optimizing the retrieval process by exploiting high-order relationships. Figure 5 illustrates the pipeline for applying GNNs in retrieval.
@@
Recent studies (Yu et al., 2022; Ye et al., 2023; Tang et al., 2024) have demonstrated that contrastive learning in GNN based retrieval significantly improves robustness. Yu et al. (2022) argue that costly graph perturbations are unnecessary; instead, generating contrastive views by injecting uniform noise into the embedding space achieves superior performance and efficiency. Subsequent research (Ye et al., 2023; Tang et al., 2024) has further enhanced model resilience by applying structure denoising and adversarial embedding perturbations. Although GNN based retrieval methods show remarkable results in understanding and leveraging complex user-item interaction patterns, enhancing their efficiency for large-scale retrieval remains a critical problem.
@@ 4.5. MLP Based Collaborative Filtering Methods
Unlike conventional collaborative filtering techniques that often rely on linear models and explicit factorization methods, neural collaborative filtering (NCF) (He et al., 2017) uses an MLP to model non‐linear user–item interactions. Concretely, let $\mathbf{e}_{u},\mathbf{e}_{i}$ be the embedding vectors for user $u$ and item $i$. NCF concatenates these embeddings as
(14) $\mathbf{h}^{(0)}_{u,i}\;=\;\bigl[\mathbf{e}_{u}\,\|\,\mathbf{e}_{i}\bigr],$
and then applies $L$ hidden layers:
(15) $\mathbf{h}^{(\ell)}_{u,i}\;=\;Activation\bigl(\mathbf{W}^{(\ell)}\,\mathbf{h}^{(\ell-1)}_{u,i}+\mathbf{b}^{(\ell)}\bigr),\quad\ell=1,\dots,L,$
where $\mathbf{W}^{(\ell)}$, $\mathbf{b}^{(\ell)}$. The final layer output $\mathbf{h}^{(L)}_{u,i}$ is then projected to a scalar score via another affine transformation and sigmoid activation.  Bai et al. (2017) enhance NCF by incorporating local neighborhood information using convolutional filters, improving interaction representation between users and items.  Liu et al. (2017) propose PhD, a hybrid recommendation model combining user-item interactions with side information using probabilistic matrix factorization.  Cao et al. (2018) develop a group recommendation method integrating an attention mechanism with MLP within the NCF framework.  Chen et al. (2018b) utilize heterogeneous information networks (HIN) to represent diverse data types, capturing complex semantics through meta paths.  Fu et al. (2018) investigate both local and global co-occurrences in user-user and item-item relations to capture contextual information more comprehensively.  Zhou et al. (2019) enrich the model’s understanding by incorporating textual representations from items, enhancing both explicit and implicit feedback.  Chen et al. (2019a) combine deep feature learning and deep interaction modeling using two MLP networks with a rating matrix.  Otunba et al. (2019) propose an ensemble framework integrating generalized matrix factorization (GMF) and MLP networks, leveraging the strengths of deep and shallow architectures.  Xue et al. (2019) focus on capturing high-order item relations in item based CF models using an NCF based approach.  Yi et al. (2019a) develop deep matrix factorization (DMF), integrating diverse side information from users and items to enhance the retrieval process.  Chen et al. (2020) propose efficient whole-data learning strategies to avoid common sampling techniques in CF retrieval methods. Additionally, closely related methods like CML (Hsieh et al., 2017) is based on metric learning, and NBPO (Yu and Qin, 2020) employs noisy-label robust learning techniques for retrieval.